OpenStudy (ellamoyseyuk):

multiply [4,0,-1,2,-3,-1]*[0,1,-3,1]

jimthompson5910 (jim_thompson5910):

use a semicolon to denote a new line so if you wanted to write out this matrix \[\Large \begin{bmatrix}3 & 0\\2 & -1\end{bmatrix}\] then you would write `[3, 0; 2, -1]`

OpenStudy (ellamoyseyuk):

oh okay

jimthompson5910 (jim_thompson5910):

So is [4,0,-1,2,-3,-1] all one row? or 2 rows?

OpenStudy (ellamoyseyuk):

3 rows

OpenStudy (ellamoyseyuk):

[4,0;-1,2;-3,-1]

jimthompson5910 (jim_thompson5910):

how about the second matrix?

OpenStudy (ellamoyseyuk):

two rows

OpenStudy (ellamoyseyuk):

[0,1;-3,1]

jimthompson5910 (jim_thompson5910):

So the first matrix is this \[\Large \begin{bmatrix}4 & 0\\-1 & 2\\-3 & -1\\ \end{bmatrix}\]

OpenStudy (ellamoyseyuk):

yes

jimthompson5910 (jim_thompson5910):

The second matrix is \[\Large \begin{bmatrix}0 & 1\\-3 & 1 \end{bmatrix}\]

OpenStudy (ellamoyseyuk):

yes

jimthompson5910 (jim_thompson5910):

We want to multiply \[\Large \begin{bmatrix}4 & 0\\-1 & 2\\-3 & -1\\ \end{bmatrix}*\begin{bmatrix}0 & 1\\-3 & 1 \end{bmatrix}\]

OpenStudy (ellamoyseyuk):

yes

jimthompson5910 (jim_thompson5910):

What steps/work have you done so far?

OpenStudy (ellamoyseyuk):

0,0,-1,2

OpenStudy (ellamoyseyuk):

i dont get it really..

jimthompson5910 (jim_thompson5910):